OpenStudy (anonymous):

Is my answer correct? What is the slope of the line 5x - 2y = -10? My answer is -5/2.

OpenStudy (raden):

false

OpenStudy (raden):

miss the negative sign there

OpenStudy (zpupster):

subtract 5x both sides -2y=-5x-10 divide -2 y=-5x/-2-10/-2 y=5/2x+5